OpenStudy (anonymous):

A circle has its center at the center of a square with 8-inch sides. Find the area of the square not covered by the circle. Round the answer to the nearest tenth. a.57.1 in2 b.9.1 in2 c.2.0 in2 d.13.7 in2

OpenStudy (anonymous):

 so the area required is, =x^2-pi(x/2)^2 =(1-pi/4)x^2 evaluate please.